what is the slope of a line that passes through the points (-4,4) and (6,-8)
Answer by ad_alta(240)

Remember that slope can be interpreted as the rate (how fast) a line changes as "x" values increase. If slope is 'm,' we can write m=(y_1-y_2)/(x_1-x_2). Thus, the slope of the line in question is (-8-4)/(6--4)=-12/10=-6/5